Required Informatlon [The following information applies to the questions displayed below] The accounting records of Nettle Distribution show the following assets and labi ities as of December 31, 2016 and 2007. Part 2 of 3 016 Cash Accounts receivable Office supplies office equipment Trucks Building Land Accounts payable Note payable $ 55,857 $1e,731 23,431 4,7143,453 144,714 154, 148 56,629 65, 629 0 188,776 47,186 38,969 135,882 0.41 points 00:40:11 78,576 eBook Print References 2. Compute net income for 2017 by comparing total equrty amounts for these two yeers and using the folowing information 2017, the owner Invested $33,000 additional cash In the business (in exchange for common stock) and the company pald $45.600 cash in dividend Equity, December 31, 2018 ... 369.570 Add: Net income 3.000 402.5701 45.600 Less: Dividends 309.5 52.600) Equity. December 31, 2017 12 Next >
Answer #1
A B C D E F G H I J
2
3 2016 2017 2016 2017
4 Cash $55,057 $10,731 Accounts Payable 78570 38969
5 Accounts Receivable $29,883 $23,431 Notes Payable 0 135882
6 Office Supplies $4,714 $3,453 Total Liabilities 78570 174851
7 Total Current Assets $89,654 $37,615
8 Office Equipment $144,714 $154,148 Stockholders' Equity $212,427 $311,423 =H12-H6
9 Trucks $56,629 $65,629
10 Building $0 $188,776
11 Land $0 $40,106
12 Total Assets $290,997 $486,274 Tota liabilities and stockholders Equity $290,997 $486,274 =E12
13
14
15 Equity Dec 31 2016 $212,427 =G8
16 Add: Net Income $111,596 =D20+D19-D15-D17
17 Additonal Cash Invested $33,000 33000
18 $357,023 =D15+D16+D17
19 Less Dividends $45,600 45600
20 Equity Dec 31 2017 $311,423 =H8
21

Formula sheet

A B C D E F G H I J
2
3 2016 2017 2016 2017
4 Cash 55057 10731 Accounts Payable 78570 38969
5 Accounts Receivable 29883 23431 Notes Payable 0 135882
6 Office Supplies 4714 3453 Total Liabilities =SUM(G4:G5) =SUM(H4:H5)
7 Total Current Assets =SUM(D4:D6) =SUM(E4:E6)
8 Office Equipment 144714 154148 Stockholders' Equity =G12-G6 =H12-H6 =getformula(H8)
9 Trucks 56629 65629
10 Building 0 188776
11 Land 0 40106
12 Total Assets =D7+SUM(D8:D11) =E7+SUM(E8:E11) Tota liabilities and stockholders Equity =D12 =E12 =getformula(H12)
13
14
15 Equity Dec 31 2016 =G8 =getformula(D15)
16 Add: Net Income =D20+D19-D15-D17 =getformula(D16)
17 Additonal Cash Invested 33000 =getformula(D17)
18 =D15+D16+D17 =getformula(D18)
19 Less Dividends 45600 =getformula(D19)
20 Equity Dec 31 2017 =H8 =getformula(D20)
21